If i am 8 and my dad is 41 then in how many years will my dad be twice my age?
Answer by Maths68(1474)

Your present age = 8
Yours Dad = 41
Let
It takes = x years
After x years
your Dad will be = 41+x
You will be = 8+x
Then
Your Dad will be twice of your age
2(8+x)=41+x
16+2x=41+x
2x-x=41-16
x=25
After 25 years your Dad will be twice of your age.
Check
=====
After 25 years you will be = 8+25=33
After 25 years your Dad will be = 41+25=66